# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v3

@@ III.1 Qualitative analysis
Note that the factor $\phi_{0}^{*}-\phi_{0}$ of the denominator coincides with the factor in $\Delta T_{c}$. The configuration with $\theta_{0}\neq 0$ and $\theta_{\rm L}\neq 0$ gradually crossovers toward one for $\bm{H}=0$. See Fig. 2(c). When $\tilde{\bm{h}}$ is sufficiently large, $u_{0}$ grows and $\delta F_{4}$ plays an important role. Thus, ${\boldsymbol{\phi}}_{\rm L}$ is locked at $\theta_{\rm L}=\pm\pi/2$ as shown in Fig. 2(d). This is because $\delta F_{4}>\delta F_{3}$ for large $\tilde{h}_{u}$. See Eq. (18). By the similar analysis to Eqs. (21)-(23), the boundary determined by the field below which $|\theta_{\rm L}|<\pi/2$ [Fig. 2(e)] is given as
@@
For sufficiently large $\lambda_{b}$, the $\theta_{\rm L}=\pm\pi/2$ phase never appears because of $\tilde{T}_{c}^{v}<0$. Of course, for extremely large $\tilde{\bm{h}}$, the L point order vanishes trivially.
Finally, we briefly comment about the cases for $u_{0}<0$ ($\theta>\theta_{111}$) under finite magnetic fields. As we have noted, there is no phase transition once a quadrupole moment with $\theta_{\rm L}=\pm\pi/2$ is ordered as shown in Fig. 2(f). This is because $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$ can minimize $\delta F_{3}$, $\delta F_{4}$, and $F_{\Gamma}$, simultaneously.
@@ III.2 Phase diagram
Now we show the results of numerical minimization of the single-$\bm{q}$ free energy (6), (10), (13), and (15). We will show the result for ${\bm{H}}\parallel[001]$, i.e., $\tilde{h}_{u}>0$ and then for ${\bm{H}}\parallel[110]$, $\tilde{h}_{u}<0$. To understand the qualitative nature of the phase diagram, it is sufficient to discuss the cases for $J_{2}=J_{3}=0$.
Figures 3(a)–3(c) show the $T$-$H$ phase diagram for $\bm{H}\parallel[001]$, where the color represents the angle for the L point moment $\theta_{\rm L}$: (a) $b=0.1$, (b) $b=0.05$ , and (c) $b=0.01$ with $c=0.5$ fixed. As discussed in the previous section, the phase III is stabilized under high magnetic fields. This phase, however, is suppressed when $b$ becomes larger. We note that the collinear phase (phase I) appears for high $T$ and intermediate $H$. The reason why the phase I cannot appear at the lower $T$ is understood by examining Eq. (23); there is no solution of the critical field $\tilde{h}_{u}^{c1}$ when $\phi_{\rm L}^{2}$, which is approximately $H$ independent, is large.
Figure 3: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, $J_{2}=J_{3}=0.0$, $J_{4}=1.0$, and $c=0.5$. (a) $b=0.1$, (b) $b=0.05$, and (c) $b=0.01$. Color map represents the L point order parameter angle $\theta_{\rm L}$.
For $\bm{H}\parallel[110]$, the situation becomes quite simpler. Figure 4(a) shows the $T$-$H$ phase diagram for $\bm{H}\parallel[110]$. As discussed in the last part in Sec. III.1, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$.
Figures 4(b) and 4(c) show the real space quadrupole configurations with the ordering vector $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively. For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on the (111) plane, while antiferroic between the neighboring (111) planes as shown in different colors in Fig. 4. When the ordering vector is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
Figure 4: (a) $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=1.0$, $J_{2}=J_{3}=0$, $J_{4}=1$, $c=0.5$, and $b=0.05$. Color map represents the L point order parameter angle $\theta_{\rm L}$. Single-$\bm{q}$ order with ${\bm{k}}={\bm{k}}_{1}$ for (b) the phase I and (c) the phase IV.
@@ IV.1 Free energy for symmetric states
${\boldsymbol{\phi}}^{\alpha}(\bm{r})=\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}=\frac{\Phi_{\alpha}}{\sqrt{\mathcal{N}}}\sum_{n=1}^{\mathcal{N}}e^{i{\bm{k}}_{n}\cdot\bm{r}}\begin{bmatrix}\cos\theta_{n}^{\alpha}\\ \sin\theta_{n}^{\alpha}\\ \end{bmatrix},$ (25)
where $\bm{r}$ is labeled by the $A$ sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same $\Phi_{\alpha}/\sqrt{\mathcal{N}}(>0)$ for simplicity (symmetric states). For $n=1$ $\bm{\phi}^{A}_{1}=\bm{\phi}^{B}_{1}$, while, for $n=2,3$ and $4$, $\bm{\phi}^{A}_{n}=-\bm{\phi}^{B}_{n}$. As mentioned before, this is due to our unit cell convention. In this normalization $(\mathcal{N})$, the quadratic part is unchanged from the single-$\bm{q}$ expression. Thus, types of order realized below the second-order transition at $T=T_{c}$ are determined by the fourth-order terms.
@@
$F_{\rm L}^{2\bm{q}(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
@@
$F_{\rm L}^{4\bm{q}(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg{[}\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}\Bigg{]}.$ (27)
@@
Note that the $c^{\prime\prime\prime}$ terms cancel when summing up the terms for $\alpha=A$ and $B$, since $\phi^{A}_{n}=(-1)^{p_{n}}\phi^{B}_{n}$. The minimum value of $F_{\rm L}^{4\bm{q}(4)}$ for $c^{\prime\prime}>0$ is realized with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=2$, where two orthogonal pairs with the relative angle between the two is arbitrary, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta^{\alpha}+\pi/2,\theta^{\alpha},\varphi^{\alpha}+\pi/2,\varphi^{\alpha})$. For $c^{\prime\prime}<0$, the configurations for the minimum free energy is, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta,\theta,\theta,\theta)$ with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=6$. The unit cell for the quadrupole-$\bm{q}$ configurations contains eight $A$-sublattice sites with sixteen sites in total. See also Fig. 9 in Appendix A.
@@
Figure 6 shows the phase diagram at $T=T_{c}$ on the $c^{\prime\prime}$$c^{\prime}$ plane for $c>0$. Note that there is no $c^{\prime\prime\prime}$ dependence. For large $c^{\prime}$, multiple-$\bm{q}$ states are not realized as is evident from Eqs. (26) and (27). The conditions given by Eq. (8) just correspond to the point $(c^{\prime}/c,c^{\prime\prime}/c)=(2,4)$, where the single-$\bm{q}$ and double-$\bm{q}$ are degenerate. As long as the local fourth-order terms are dominant, quadrupole-$\bm{q}$ orders are unlikely. Thus, the $c^{\prime}$ term is important when one compares the free energy for the single- and double-$\bm{q}$ orders.
@@
Figure 5: Phase diagram at $T=T_{c}$ in the $c^{\prime}$$c^{\prime\prime}$ plane for $c>0$. Filled circle indicate the bare parameter set $c^{\prime}=2c$ and $c^{\prime\prime}=4c$ given by Eq. (8). Region for $c^{\prime}+c^{\prime\prime}<-2c$ is unphysical since the fourth-order term becomes negative. For quadruple-$\bm{q}$ states, their representative configurations for $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})$ are indicated.
@@
So far, we have treated the fourth-order coefficients $c$’s as a phenomenological parameters. Here, we discuss the renormalization owing to modes other than $\bm{\phi}_{0}$ and $\bm{\phi}_{n}$. In the lowest orders, the cubic terms in Eq. (5) are the most important. Among them, there are special terms which connect two L-point modes at $\bm{q}_{m}$ and $\bm{q}_{n}$ ($m\neq n$) via $\sim\phi_{\bm{k}}\phi_{m}\phi_{n}$, where ${\bm{k}}$ is one of the wavenumbers at the X points $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, or $\bm{q}_{\rm Z}=(0,0,2\pi)$. The combination of $({\bm{k}},\bm{q}_{m},\bm{q}_{n})$ is restricted by the Umklapp condition ${\bm{k}}+\bm{q}_{m}+\bm{q}_{n}\equiv\bf{0}$ and there are two sets of $(m,n)$ for a given ${\bm{k}}$. As shown in Appendix B, these couplings lead to suppression of the coefficients $c^{\prime}$ and $c^{\prime\prime\prime}$. See Eq. (38). Furthermore, the fifth-order couplings between the $\Gamma$ and the L points fields appear after integrating out the X point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$. They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$, and thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
@@
The suppression of $c^{\prime}$ and $c^{\prime\prime\prime}$ due to the renormalization with retaining $c^{\prime\prime}>0$ suggests that the orthogonal double-$\bm{q}$ states are more stable. However, the coupling between the fields at the $\Gamma$ and the L points possesses an energy gain for the single-$\bm{q}$ orders and double-$\bm{q}$ ones with $\phi_{m}\neq\phi_{n}$. This is because there is a finite $\phi_{0}$ induced in the single-$\bm{q}$ states, while there is no gain for the double-$\bm{q}$ states with $\phi_{n}=\phi_{m}$ and $\theta_{n}-\theta_{m}=\pm\pi/2$. See the coupling in Eq. (13). As we will see in Sec. IV.2, the fifth-order terms generated by the integration of the X points fields choose some particular $\theta$ among degenerate orthogonal double-$\bm{q}$ configurations.
@@ IV.2 Phase diagram
Now, we discuss $T$$H$ phase diagram ($H\equiv|\bm{H}|$) taking into account multiple-$\bm{q}$ orders in addition to the single-$\bm{q}$ ones. We numerically minimize the free energy $F_{\rm L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}+\delta F$, where $\delta F$ is due to the integration of the X point fields as discussed in Appendix B and the others are given in Eqs. (6), (10), (13), and (15). For simplify the analysis, we assume ${\boldsymbol{\phi}}_{n}^{B}={\boldsymbol{\phi}}_{n}^{A}(-1)^{p_{n}}$ valid for $J_{\rm L}<0$ and do not take into account quadruple-$\bm{q}$ orders, since their energy is much higher than single- and double-$\bm{q}$ states for the conventional fourth-order terms [Eq. (8)] as shown in Fig. 6. Figures 6 and 7 show phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, and those for $\bm{H}\parallel[110]$ with $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\leq 0)$, respectively. The representative double-$\bm{q}$ configurations in the quadrupole space and real space are also illustrated in Fig. 8.
@@
Figure 6: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$. (a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$. Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, where double-$\bm{q}$ orders with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ are assumed. For single-$\bm{q}$ orders, we use the convention of $\phi_{2}=0$. Thick lines represent phase boundaries, where “1st” represents a line of first-order phase transition. Dashed lines are boundaries among double-$\bm{q}$ configurations. Low $T$ regime is not shown since the Landau expansion is not valid in general.
@@
Figure 6(a) shows the phase diagram for $J_{1}=1$, $c=0.5$, and $b=0.1$. The exchange coupling constants are the same as used in Fig. 3: $(J_{2},J_{3},J_{4})=(0,0,1)$. Except for the low-field region, the phase diagram is similar to those shown in Fig. 3, showing the single-$\bm{q}$ orders I, II, and III. An orthogonal double-$\bm{q}$ configuration I2q with $\bm{\phi}_{{\bm{k}}_{1}}=(\phi_{1},0)$ and $\bm{\phi}_{{\bm{k}}_{2}}=(0,\phi_{2})$ is stabilized for finite magnetic fields at high $T$ region below the phase I. Here, we take a domain with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. The configuration of I2q state in the quadrupole space is shown in Fig. 8(a) and its real-space one is illustrated in Fig. 8(b). At ${\bm{H}}=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec. IV.1 are realized. As $H$ increases, $\phi_{1}$ becomes larger than $\phi_{2}$. Since the phase I is the single-$\bm{q}$ order with $v_{\rm L}=0$, the symmetry breaking from the phase I to the phase I2q is caused by the emergence of $\bm{\phi}_{2}=(0,\phi_{2})$ through the second-order transition. In this sense, $\phi_{1}>\phi_{2}$ is quite natural. In the free energy, this is due to the $\delta F_{3}$, which is minimized when the L point field is parallel to $\bm{\phi}_{0}=(u_{0},0)$, while maximized when it is perpendicular. As for the phase boundary between the phase I2q and the phase II, it is of first-order, while the other phase boundaries are second-order.

### v4

@@ III.1 Qualitative analysis
Note that the factor $\phi_{0}^{*}-\phi_{0}$ of the denominator coincides with the factor in $\Delta T_{c}$. The configuration with $\theta_{0}\neq 0$ and $\theta_{\rm L}\neq 0$ gradually crosses over to the configuration for $\bm{H}=0$. See Fig. 3(c). When $\tilde{\bm{h}}$ is sufficiently large, $u_{0}$ grows and $\delta{\mathcal{F}}_{4}$ plays an important role. Thus, ${\bm{\phi}}_{\rm L}$ is locked at $\theta_{\rm L}=\pm\pi/2$ as shown in Fig. 3(d). This is because $\delta{\mathcal{F}}_{4}>\delta{\mathcal{F}}_{3}$ for large $\tilde{h}_{u}$. See Eq. (18). By the similar analysis to Eqs. (21)–(23), the boundary determined by the field below which $|\theta_{\rm L}|<\pi/2$ [Fig. 3(e)] is given as
@@
For sufficiently large $\lambda_{b}$, the $\theta_{\rm L}=\pm\pi/2$ phase never appears because of $\tilde{T}_{c}^{v}<0$. Of course, for extremely large $\tilde{\bm{h}}$, the L-point order vanishes trivially.
Finally, we briefly comment about the cases for $u_{0}<0$ ($\Theta>\Theta_{111}$) under finite magnetic fields. As we have noted, there is no phase transition once a quadrupole moment with $\theta_{\rm L}=\pm\pi/2$ is ordered as shown in Fig. 3(f). This is because $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$ can minimize $\delta{\mathcal{F}}_{3}$, $\delta{\mathcal{F}}_{4}$, and ${\mathcal{F}}_{\Gamma}$, simultaneously.
@@ III.2 Phase diagram
Now we show the results of numerical minimization of the single-$\bm{q}$ free energy (6), (10), (13), and (15). We will discuss the result for ${\bm{H}}\parallel[001]$, i.e., $\tilde{h}_{u}>0$ and then for ${\bm{H}}\parallel[110]$, $\tilde{h}_{u}<0$. To understand the qualitative nature of the phase diagram, it is sufficient to discuss the cases for $J_{2}=J_{3}=0$.
Figures 4(a)–4(c) show the TTHH phase diagram for $\bm{H}\parallel[001]$, where $H=|\bm{H}|$ and the color represents the angle for the L point moment $\theta_{\rm L}$: (a) $b=0.1$, (b) $b=0.05$, and (c) $b=0.01$ with $c=0.5$ fixed. As discussed in the previous section, the phase III is stabilized under high magnetic fields. This phase, however, is suppressed when $b$ becomes larger. We note that the collinear phase (phase I) appears for high $T$ and intermediate $H$. The reason why the phase I cannot appear at the lower $T$ is understood by examining Eq. (23); there is no solution of the critical field $\tilde{h}_{u}^{c1}$ when $\phi_{\rm L}^{2}$, which is approximately $H$ independent, is large.
For $\bm{H}\parallel[110]$, the situation becomes much simpler. Figure 4(d) shows the TTHH phase diagram for $\bm{H}\parallel[110]$. As discussed in the last part in Sec. III.1, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$. For $b=0.1$ and $b=0.01$, the phase diagrams are similar and we do not show them.
Figures 4(e) and 4(f) show the real space quadrupole configurations with the ordering wave number $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively. For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on the (111) planes, while antiferroic between the neighboring (111) planes as shown in Fig. 4. Namely the quadrupole moments at the sites indicated by white spheres and those by red ones have opposite sign. When the ordering wave number is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
@@ IV.1 Free energy for symmetric states
${\bm{\phi}}^{\alpha}(\bm{r})=\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}=\frac{\Phi_{\alpha}}{\sqrt{\mathcal{N}}}\sum_{n=1}^{\mathcal{N}}e^{i{\bm{k}}_{n}\cdot\bm{r}}\begin{bmatrix}\cos\theta_{n}^{\alpha}\\ \sin\theta_{n}^{\alpha}\\ \end{bmatrix}$, (25)
where $\bm{r}$ is labeled by the $A$-sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same $\Phi_{\alpha}/\sqrt{\mathcal{N}}~(>0)$ for simplicity (symmetric states). For $n=1$ $\bm{\phi}^{A}_{1}=\bm{\phi}^{B}_{1}$, while, for $n=2,3$ and $4$, $\bm{\phi}^{A}_{n}=-\bm{\phi}^{B}_{n}$. As mentioned before, this is due to our unit cell convention. In this normalization $(\mathcal{N})$, the quadratic part is unchanged from the single-$\bm{q}$ expression. Thus, types of order realized below the second-order transition at $T=T_{c}$ are determined by the fourth-order terms.
@@
${\mathcal{F}}_{\rm L}^{2\bm{q}(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
@@
${\mathcal{F}}_{\rm L}^{4\bm{q}(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg[\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}\Bigg].$ (27)
@@
Note that the $c^{\prime\prime\prime}$ terms cancel when summing up the terms for $\alpha=A$ and $B$, since $\phi^{A}_{n}=(-1)^{p_{n}}\phi^{B}_{n}$. The minimum value of ${\mathcal{F}}_{\rm L}^{4\bm{q}(4)}$ for $c^{\prime\prime}>0$ is realized with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=2$, where two orthogonal pairs with the relative angle between the two is arbitrary, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta^{\alpha}+\pi/2,\theta^{\alpha},\varphi^{\alpha}+\pi/2,\varphi^{\alpha})$. For $c^{\prime\prime}<0$, the configurations for the minimum free energy is, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta,\theta,\theta,\theta)$ with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=6$. The unit cell for the quadrupole-$\bm{q}$ configurations contains eight $A$-sublattice sites with sixteen sites in total. See also Fig. 9 in Appendix A.
@@
Figure 5 shows the phase diagram at $T=T_{c}$ on the $^{\prime\prime}cc^{\prime}$ plane for $c>0$. Note that there is no $c^{\prime\prime\prime}$ dependence. For large $c^{\prime}$, multiple-$\bm{q}$ states are not realized as is evident from Eqs. (26) and (27). The conditions given by Eq. (8) just correspond to the point $(c^{\prime}/c,c^{\prime\prime}/c)=(2,4)$, where the single-$\bm{q}$ and double-$\bm{q}$ orders are degenerate. As long as the local fourth-order terms are dominant, quadruple-$\bm{q}$ orders are unlikely. Thus, the $c^{\prime}$ term is important when one compares the free energy for the single- and double-$\bm{q}$ orders.
@@
Figure 5: Phase diagram at $T=T_{c}$ in the $^{\prime}c′′c^{\prime\prime}$ plane for $c>0$. A filled circle indicates the bare parameter set $c^{\prime}=2c$ and $c^{\prime\prime}=4c$ given by Eq. (8). Region for $c^{\prime}+c^{\prime\prime}<-2c$ is unphysical since the fourth-order term becomes negative. For quadruple-$\bm{q}$ states, their representative configurations for $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})$ are indicated.
@@
So far, we have treated the fourth-order coefficients $c$’s as phenomenological parameters. Here, we discuss the renormalization owing to modes other than $\bm{\phi}_{0}$ and $\bm{\phi}_{n}$. In the lowest orders, the cubic terms in Eq. (5) are the most important. Among them, there are special terms which connect two L-point modes at $\bm{q}_{m}$ and $\bm{q}_{n}$ ($m\neq n$) via $\sim\phi_{\bm{k}}\phi_{m}\phi_{n}$, where $\bm{k}$} is one of the wavenumbers at the X points $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, or $\bm{q}_{\rm Z}=(0,0,2\pi)$. The combination of $({\bm{k}},\bm{q}_{m},\bm{q}_{n})$ is restricted by the Umklapp condition ${\bm{k}}+\bm{q}_{m}+\bm{q}_{n}\equiv\bf{0}$ and there are two sets of $(m,n)$ for a given $\bm{k}$}. As shown in Appendix B, these couplings lead to suppression of the coefficients $c^{\prime}$ and $c^{\prime\prime\prime}$. See Eq. (38). Furthermore, the fifth-order couplings between the $\Gamma$- and the L-points fields appear after integrating out the X-point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$. They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$. Thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
@@
The suppression of $c^{\prime}$ and $c^{\prime\prime\prime}$ due to the renormalization with retaining $c^{\prime\prime}>0$ suggests that the orthogonal double-$\bm{q}$ states are more stable. However, the coupling between the fields at the $\Gamma$ and the L points possesses an energy gain for the single-$\bm{q}$ orders and double-$\bm{q}$ ones with $\phi_{m}\neq\phi_{n}$. This is because there is a finite $\phi_{0}$ induced in the single-$\bm{q}$ states, while there is no gain for the double-$\bm{q}$ states with $\phi_{n}=\phi_{m}$ and $\theta_{n}-\theta_{m}=\pm\pi/2$. See the coupling in Eq. (13). As we will see in Sec. IV.2, the fifth-order terms generated by the integration of the X-points fields choose some particular $\theta_{n}$’s among degenerate orthogonal double-$\bm{q}$ configurations.
@@ IV.2 Phase diagram
Now, we discuss TTHH phase diagram ($H=|\bm{H}|$) taking into account multiple-$\bm{q}$ orders in addition to the single-$\bm{q}$ ones. We numerically minimize the free energy ${\mathcal{F}}_{\rm L}+{\mathcal{F}}_{\Gamma}+\delta{\mathcal{F}}_{3}+\delta{\mathcal{F}}_{4}+\delta{\mathcal{F}}_{\rm{XL}}$, where $\delta{\mathcal{F}}_{\rm{XL}}$ is due to the integration of the X-point fields as discussed in Appendix B and the others are given in Eqs. (6), (10), (13), and (15). To simplify the analysis, we assume ${\bm{\phi}}_{n}^{B}={\bm{\phi}}_{n}^{A}(-1)^{p_{n}}$ valid for $J_{\rm L}<0$ and do not take into account quadruple-$\bm{q}$ orders, since their energy is much higher than single- and double-$\bm{q}$ states for the conventional fourth-order terms [Eq. (8)] as shown in Fig. 5. Figures 6(a)–6(d) show the phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, $(J_{2},J_{3},J_{4})=(0.0,0.0,1.0)$ with (a) $b=0.1$ and (b) $b=0.05$. For $\bm{H}\parallel[110]$ with $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\leq 0)$, the corresponding phase diagrams are shown in Figs. 6(c) and 6(d). One can notice that there are two double-$\bm{q}$ phases labeled by I2q and II2q, in addition to the single-$\bm{q}$ phases I, II, III, and IV as already discussed above.
@@
Figure 6: TTHH phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$. (a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$. Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, where double-$\bm{q}$ orders with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ are assumed. For single-$\bm{q}$ orders, we use the convention of $\phi_{2}=0$. Thick lines represent phase boundaries, where “1st” represents a line of first-order phase transition. Dashed lines are boundaries among double-$\bm{q}$ configurations. Low $T$ regime is not shown since the Landau expansion is not valid in general.
@@
Figure 6(a) shows the phase diagram for $J_{1}=1$, $c=0.5$, and $b=0.1$. The exchange coupling constants are the same as used in Fig. 4: $(J_{2},J_{3},J_{4})=(0.0,0.0,1.0)$. Except for the low-field region, the phase diagram is similar to those shown in Fig. 4, showing the single-$\bm{q}$ orders I, II, and III. An orthogonal double-$\bm{q}$ configuration I2q with $\bm{\phi}_{{\bm{k}}_{1}}=(\phi_{1},0)$ and $\bm{\phi}_{{\bm{k}}_{2}}=(0,\phi_{2})$ is stabilized for finite magnetic fields, touching its high-$T$ side boundary with the phase I. The configuration of I2q state in the quadrupole space is shown in Fig. 7(a) and its real-space one is illustrated in Fig. 7(b), where a domain with the ordering wave numbers ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ is shown. At $H=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec. IV.1 are realized. As $H$ increases, $\phi_{1}$ becomes larger than $\phi_{2}$. Since the phase I is the single-$\bm{q}$ order with $v_{\rm L}=0$, the symmetry breaking from the phase I to the phase I2q is caused by the emergence of $\bm{\phi}_{2}=(0,\phi_{2})$ through the second-order transition. In this sense, $\phi_{1}>\phi_{2}$ is quite natural. In the free energy, this is due to the $\delta{\mathcal{F}}_{3}$, which is minimized when the L point field is parallel to $\bm{\phi}_{0}=(u_{0},0)$, while maximized when it is perpendicular. As for the phase boundary between the phase I2q and the phase II, it is of first-order, while the other phase boundaries are second-order.
[…]